# Investigation of the Impact of Synthetic Training Data in the Industrial Application of Terminal Strip Object Detection

### v1

@@ Abstract
In industrial manufacturing, numerous tasks of visually inspecting or detecting specific objects exist that are currently performed manually or by classical image processing methods. Therefore, introducing recent deep learning models to industrial environments holds the potential to increase productivity and enable new applications. However, gathering and labeling sufficient data is often intractable, complicating the implementation of such projects. Hence, image synthesis methods are commonly used to generate synthetic training data from 3D models and annotate them automatically, although it results in a sim-to-real domain gap. In this paper, we investigate the sim-to-real generalization performance of standard object detectors on the complex industrial application of terminal strip object detection. Combining domain randomization and domain knowledge, we created an image synthesis pipeline for automatically generating the training data. Moreover, we manually annotated 300 real images of terminal strips for the evaluation. The results show the cruciality of the objects of interest to have the same scale in either domain. Nevertheless, under optimized scaling conditions, the sim-to-real performance difference in mean average precision amounts to 2.69% for RetinaNet and 0.98% for Faster R-CNN, qualifying this approach for industrial requirements.
@@ 1 INTRODUCTION
Currently, various research fields in machine learning (ML) are evolving rapidly, setting new benchmarks and demonstrating the potential benefit of this technology for society. In particular, the availability of computing power and large-scale public datasets advanced the development and initiated the era of deep learning (DL). However, the amount of data required to train DL models limits the transferability of current state-of-the-art approaches to real-world applications in which data is scarce and difficult to gather [Tsirikoglou et al., 2020], such as computer vision (CV) applications in industrial manufacturing.
The main challenge arises from the manual effort to collect and label the training images for supervised learning. For instance, the creators of Microsoft Common Objects in Context (MS COCO), a commonly used benchmark dataset for object detection and semantic segmentation, stated in [Lin et al., 2014] that it took around 60,000 worker hours to gather and annotate these images, which is intractable for most industrial applications. Moreover, industrial settings are dynamic, so the environment and the objects of interest may change, forcing the training data to be dynamically adaptable within a reasonable time. Therefore, recent works focused on the usage of simulated images for model training and investigated their impact on model performance in real-world test cases [Akar et al., 2022, Ljungqvist et al., 2023, Zhu et al., 2023].
@@
Apart from the usual data augmentation techniques and the related cut-and-paste approach [Dvornik et al., 2018], domain adaptation and domain randomization (DR) are the two main concepts currently used in this area based on rendered images of 3D models. The former maps one domain into the other or both into an intermediate third domain, using, e.g., GANs [Rojtberg et al., 2020, Shrivastava et al., 2017, Wang et al., 2022]. Thus, domain adaptation still relies on real images, which might be unsuitable for industrial use cases. In contrast, DR is based on random, even unrealistic variations, e.g., of textures, backgrounds, or lighting conditions, to suggest to the DL model that real images are just another variation of the learned domain [Tobin et al., 2017]. Hence, it aims to generalize directly to the real world and enables fully synthetic training, which has already shown promising results in different use cases, such as detecting cars [Tremblay et al., 2018], small load carriers [Mayershofer et al., 2021], and turbine blades at a manual working station [Eversberg and Lambrecht, 2021]. Nevertheless, experiments in industrial environments were conducted under highly simplified conditions, including a small amount of well-separable classes and considering only similar-sized objects. The expectable performance of state-of-the-art approaches on more complex industrial tasks remains unclear.
@@
We approached this task by building an automated image synthesis pipeline and generating 30,000 synthetic images of random terminal strips. Moreover, we annotated 300 handmade real images for evaluation and provide baseline results of two standard object detectors, entirely trained on synthetic images and tested on real images. As we found that scaling is crucial in this specific use case, we also show how synthetic data can be used to learn a suitable image preprocessing that significantly improves the detection performance. To motivate further developments on this use case, we made our dataset publicly available on Zenodo [^2] https://zenodo.org/records/10674995 for research purposes.
The remainder of this paper is structured as follows. Section 2 points out usual image synthesis methods and available synthetic image datasets to show recent developments in sim-to-real DL. Then, we describe in Section 3 how DR and domain knowledge are combined in our pipeline to create a synthetic image dataset for terminal strip object detection. While we detail our experiments to assess the sim-to-real domain gap and present the corresponding results in Section 4, Section 5 discusses the significance of the results for industrial manufacturers. The paper is concluded in Section 6 also giving a brief outlook for future work.
@@ 2.1 Synthetic Image Datasets
Although annotated synthetic image datasets can be generated much faster than their real-world counterparts, they are still outnumbered and only available for specific use cases. One very active research area dealing with simulated training data is autonomous driving, in which different approaches are used for dataset creation. In [Richter et al., 2016], e.g., the authors exploited the video game engine of GTA5 to obtain 25,000 photorealistic street scene images with pixel-level segmentation masks. By substituting $\nicefrac{{2}}{{3}}$ of the real training images, they could improve the performance of a semantic segmentation model, which demonstrates the potential benefit of synthetic data.
@@
Contrarily, the creators of Virtual KITTI semi-automatically reconstructed real-world street scenes from the original KITTI dataset [Geiger et al., 2012] and varied several parameters, such as the weather conditions, resulting in 35 synthetic videos with approximately 17,000 frames labeled for various video understanding tasks [Gaidon et al., 2016]. Their results show that a combination of synthetic pre-training and real fine-tuning yields better performance than each separate approach in multi-object tracking, supporting the importance of simulated data.
@@
Another notable street scene dataset is Synscapes [Wrenninge and Unger, 2018] which represents the synthetic counterpart to Cityscapes [Cordts et al., 2016]. In contrast to the previously mentioned datasets, Synscapes was created using a parameter-based random generator to build a new scene for each image. Furthermore, the developers leveraged photorealistic rendering techniques and found that high image realism is necessary to bridge the sim-to-real domain gap [Wrenninge and Unger, 2018]. Nevertheless, the application of autonomous driving is out of scope for industrial manufacturers, so the observed results may not apply to their use cases.
@@
The amount of synthetic image datasets for industrial requirements is small, even though several approaches for image synthesis have been proposed and evaluated in small-scale industrial test cases, as discussed in the following subsection. Two recently published datasets are SIP-17 [Zhu et al., 2023] and SORDI [Akar et al., 2022], aiming to close this gap. SIP-17 is a synthetic dataset for industrial parts classification consisting of 17 objects in six different use cases to allow research in close-to-reality scenarios. The experiments reveal that classification is difficult when classes share a similar albedo or have a comparable shape [Zhu et al., 2023], which is typical for industrial objects.
@@
While SIP-17 contains 33,000 images, SORDI was presented as the first large-scale synthetic image dataset for the industry with 200,000 images at the publication date. It is characterized by eight industrial assets in 32 fully synthetic factory scenarios and enables the sim-to-real benchmarking of vision models for intelligent robots in factory environments [Akar et al., 2022]. However, SORDI cannot serve the entire range of industrial applications, so further work is required to cover, e.g., quality control in assembly and object tracking.
@@
To conclude, it is difficult for industrial manufacturers to estimate how a synthetically trained DL model would perform in a specific use case, as there is not enough data available to benchmark sim-to-real methods in different industry-related applications. Thus, we provide our dataset to enable further investigations in this area with a focus on narrow industrial objects that are densely arranged and partly occlude each other.
@@ 2.2 Image Synthesis for Deep Learning
Generating synthetic training images for supervised learning tasks requires suitable design choices to improve the sim-to-real generalization performance. These choices include textures of the objects of interest, the image background, different viewpoints, lighting conditions, etc., potentially making the modeling of virtual scenes very time-consuming and complex as well. Especially when the reality should be closely approximated, modeling the entire range of possible real-world variations might eliminate the advantage of saving time for data annotation. Hence, there are various approaches to apply image synthesis that attempt to minimize the domain gap while avoiding excessive modeling effort. For instance, the authors of [Dwibedi et al., 2017] used segmentation masks of real images to cut out the objects of interest and place them onto random background images, which can be seen as image synthesis. This idea was later improved in [Dvornik et al., 2018] proposing a context model that provides patches of the background images likely to contain certain objects to create more realistic training data. Although this method is very efficient and showed promising results in object detection [Dvornik et al., 2018], it relies on real-world data which might not be available or too expensive to gather in industrial applications. Alternatively, the appearance of real-world objects can be approximated utilizing accurate 3D models and modern rendering techniques such as in [Su et al., 2015]. Here, the authors overlaid background images with object renderings, similar to the cut-and-paste approach but completely removing the necessity of annotated real-world data.
@@
Since real images are highly affected by the environment in which they are taken, the appearance of objects changes with their surroundings, impeding model prediction. Hence, a common approach is to randomize the environment of the objects of interest during image synthesis to train the DL model to deal with these variations. However, the authors of [Tobin et al., 2017] hypothesized that complete DR, including even unrealistic variations of the objects of interest, helps to close the sim-to-real domain gap as the DL model is forced to focus on the object’s shape rather than its color or glossiness. Therefore, they also randomized the textures of the 3D models and obtained promising results in the localization of simple geometric objects, enabling a robot to grab these objects in a real-world test [Tobin et al., 2017].
@@
Based on this idea, several experiments have been conducted, assessing different degrees of DR and how domain knowledge can be used during simulation to improve real-world performance. For instance, [Prakash et al., 2019] introduced context splines to the synthetic scenes to ensure that cars are always positioned on streets and [Mayershofer et al., 2021] proposed to define object relations via configuration files to set the objects of interest into more realistic contexts. However, these methods were evaluated on small individual test cases, so the results are hardly comparable and also partly contradicting. This was addressed by the authors of [Eversberg and Lambrecht, 2021] who compared numerous DR approaches in the application of detecting turbine blades at manual working stations to provide a guideline on designing image synthesis pipelines for industrial use cases. They observed it to be most beneficial to focus on realistic lighting conditions and object textures while the image background and distractors can be randomized completely [Eversberg and Lambrecht, 2021].
@@
Complementarily, in our research, we follow these guidelines and make a few application-specific changes to demonstrate the sim-to-real performance of state-of-the-art image synthesis techniques in the challenging industrial application of terminal strip object detection.
@@ 3 IMAGE SYNTHESIS PIPELINE
  • The most informative view of a terminal strip is the front view, as illustrated in Figure 1.
Since it has been shown to be beneficial to include domain knowledge in DR [Prakash et al., 2019, Mayershofer et al., 2021], we use this information in our pipeline along with standard randomization techniques to obtain random synthetic images that are contextually close to reality. In this section, we provide details about the different generation steps we implemented using the open-source rendering software Blender [^3] Blender 3.3: https://www.blender.org.
@@ 3.1 Creation of Terminal Strips
Apart from wiring, the main distractors of terminal strips are their additional accessories, such as plug-in bridges, test adapters, and markings. Although the authors of [Eversberg and Lambrecht, 2021] found that realistic distractors do not necessarily improve the performance of DL models, we decided to stay with realism for two reasons. First, the modeling effort is almost identical to simple, unrealistic distractors, as the required 3D models are available, and the virtual materials are similar to the materials used for the terminal blocks. Second, some applications may focus on detecting these accessories, e.g., to determine whether plug-in bridges are plugged correctly. Thus, each terminal strip is randomly equipped with markings, plug-in bridges, and test adapters according to its configuration.
@@ 3.3 Lighting Conditions
In recent image synthesis methods for visual DL applications, there are two common approaches to randomize the lighting conditions: using parameterizable, virtual light sources and implementing image-based lighting. The former is very flexible and allows accurate modeling of the desired illumination. However, creating real-world conditions is laborious since light rays of different colors and intensities might enter the scene from various directions, and reflections should also be considered. In contrast, image-based lighting maps a $360^{\circ}$ image on a globe and uses the pixel values to illuminate the 3D models as if they were in the captured environment. This is less customizable but reduces the creation of complex lighting conditions to the selection of suitable images and has shown to be superior when only little effort is invested in parameterizing virtual light sources [Eversberg and Lambrecht, 2021].
@@ 3.4 Image Background
Similar to the lighting conditions, synthetic image backgrounds can be virtually created and rendered along with the objects of interest [Akar et al., 2022] or they rely on real images [Eversberg and Lambrecht, 2021]. On the one hand, modeling the entire background leads to more realistic images, as all objects can be related, and visualizing realistic shadows is possible. On the other hand, random image-based backgrounds have shown promising results [Eversberg and Lambrecht, 2021] and are less complex, which is more suitable for industrial manufacturers. Nevertheless, the authors of [Alghonaim and Johns, 2021] found that realistic shadows significantly improve the performance of DL models in a 6D pose estimation task. Therefore, we decided to follow a compromising strategy, reusing the HDRIs of the environment lighting and visualizing the shadow of the terminal strip as if it were mounted on a flat surface.
@@
The implementation is based on a surrounding semi-ellipsoid, as illustrated in Figure 4, that serves as a shadow catcher. It is positioned so that the flat side represents the mounting surface for the DIN rail to which the terminal strip is centered, allowing the objects of interest to cast shadows on the background. To insert the HDRI, the same image from the environment lighting is mapped onto the surface of the semi-ellipsoid, and the respective rotation around the vertical axis is applied. As a result, the background of a synthetic image consists of the ground of the environment captured by the HDRI with additional shadows cast by the terminal strip, which is more realistic than using fully randomized backgrounds and less expensive than modeling the entire scene. See the appendix for synthetic sample images generated using this approach.

### v2

@@ Abstract
In industrial manufacturing, deploying deep learning models for visual inspection is mostly hindered by the high and often intractable cost of collecting and annotating large-scale training datasets. While image synthesis from 3D CAD models is a common solution, the individual techniques of domain and rendering randomization to create rich synthetic training datasets have been well studied mainly in simple domains. Hence, their effectiveness on complex industrial tasks with densely arranged and similar objects remains unclear. In this paper, we investigate the sim-to-real generalization performance of standard object detectors on the complex industrial application of terminal strip object detection, carefully combining randomization and domain knowledge. We describe step-by-step the creation of our image synthesis pipeline that achieves high realism with minimal implementation effort and explain how this approach could be transferred to other industrial settings. Moreover, we created a dataset comprising 30.000 synthetic images and 300 manually annotated real images of terminal strips, which is publicly available for reference and future research. To provide a baseline as a lower bound of the expectable performance in these challenging industrial parts detection tasks, we show the sim-to-real generalization performance of standard object detectors on our dataset based on a fully synthetic training. While all considered models behave similarly, the transformer-based DINO model achieves the best score with 98.40% mean average precision on the real test set, demonstrating that our pipeline enables high quality detections in complex industrial environments from existing CAD data and with a manageable image synthesis effort.
keywords:
Object Detection, Image Synthesis, Domain Randomization, Domain Gap, Terminal Strip
@@ 1 Introduction
Currently, various research fields in machine learning (ML) are evolving rapidly, setting new benchmarks and demonstrating the potential benefit of this technology for society. In particular, the availability of computing power and large-scale public datasets advanced the development and initiated the era of deep learning (DL). However, the amount of data required to train DL models limits the transferability of current state-of-the-art approaches to real-world applications in which data is scarce and difficult to gather [Tsirikoglou.2020], such as computer vision (CV) applications in industrial manufacturing.
The main challenge arises from the manual effort to collect and label the training images for supervised learning. For instance, the creators of Microsoft Common Objects in Context (MS COCO), a commonly used benchmark dataset for object detection and semantic segmentation, stated in [Lin.2014] that it took around 60,000 worker hours to gather and annotate these images, which is intractable for most industrial applications. Moreover, industrial settings are dynamic, so the environment and the objects of interest may change, forcing the training data to be dynamically adaptable within a reasonable time. Therefore, recent works focused on the usage of simulated images for model training and investigated their impact on model performance in real-world test cases [Akar.2022, Ljungqvist.2023, Zhu.2023].
@@
Apart from the usual data augmentation techniques and the related cut-and-paste approach [Dvornik.2018], domain adaptation and domain randomization (DR) are the two main concepts currently used in this area based on rendered images of 3D models. The former maps one domain into the other or both into an intermediate third domain, using, e.g., GANs [Rojtberg.2020, Shrivastava.2017, Wang.2022]. Thus, domain adaptation still relies on real images, which might be unsuitable for industrial use cases. In contrast, DR is based on random, even unrealistic variations, e.g., of textures, backgrounds, or lighting conditions, to suggest to the DL model that real images are just another variation of the learned domain [Tobin.2017]. Hence, it aims to generalize directly to the real world and enables fully synthetic training, which has already shown promising results in different use cases, such as detecting cars [Tremblay.2018], small load carriers [Mayershofer.2021], and turbine blades at a manual working station [Eversberg.2021]. Nevertheless, experiments in industrial environments were conducted under highly simplified conditions, including a small amount of well-separable classes and considering only similar-sized objects. The expectable performance of state-of-the-art approaches on more complex industrial tasks in which the objects of interest could be very similar and densely arranged remains unclear.
@@
We approached this task by building an automated image synthesis pipeline and generating 30,000 synthetic images of random terminal strips. Thereby, we focused on including domain specific characteristics, e.g., the arrangement of the objects of interest relative to each other and their colors, while randomizing the lighting conditions, viewpoints, and background to keep the modeling effort manageable. This procedure could also be transferred to other industrial applications, as we discuss in Section 5. Moreover, we annotated 300 handmade real images for evaluation and provide baseline results of four standard object detectors, without excessive hyperparameter tuning and entirely trained on synthetic images. This serves as a lower bound of the expectable sim-to-real generalization performance since any labeled ground truth data of the target domain would presumably lead to an improvement.
As we found that scaling is most crucial for the sim-to-real domain gap in this specific use case, we also show how synthetic data can be used to learn a dynamic image preprocessing that significantly improves the real-world detection performance. To motivate further developments on this use case, we made our dataset publicly available on Zenodo [^2] https://zenodo.org/records/10674995 for research purposes.
The remainder of this paper is structured as follows. Section 2 points out usual image synthesis methods and available synthetic image datasets to show recent developments in sim-to-real DL. Then, we describe in Section 3 how DR and domain knowledge are combined in our pipeline to create a synthetic image dataset for terminal strip object detection. While we detail our experiments to assess the sim-to-real domain gap and present the corresponding results in Section 4, Section 5 discusses the significance of the results for industrial manufacturers, the transferability of our strategy to other industrial scenarios, and how synthetic data performs against augmented data in the worst case consideration of no labeled ground truth images of the target domain being available. The paper is concluded in Section 6 also giving a brief outlook for future work.
@@ 2.1 Synthetic Image Datasets
Although annotated synthetic image datasets can be generated much faster than their real-world counterparts, they are still outnumbered and only available for specific use cases. One very active research area dealing with simulated training data is autonomous driving, in which different approaches are used for dataset creation. In [Richter.2016], e.g., the authors exploited the video game engine of GTA5 to obtain 25,000 photorealistic street scene images with pixel-level segmentation masks. By substituting $\nicefrac{{2}}{{3}}$ of the real training images, they could improve the performance of a semantic segmentation model, which demonstrates the potential benefit of synthetic data.
@@
Contrarily, the creators of Virtual KITTI semi-automatically reconstructed real-world street scenes from the original KITTI dataset [Geiger.2012] and varied several parameters, such as the weather conditions, resulting in 35 synthetic videos with approximately 17,000 frames labeled for various video understanding tasks [Gaidon.2016]. Their results show that a combination of synthetic pre-training and real fine-tuning yields better performance than each separate approach in multi-object tracking, supporting the importance of simulated data.
@@
Another notable street scene dataset is Synscapes [Wrenninge.2018] which represents the synthetic counterpart to Cityscapes [Cordts.2016]. In contrast to the previously mentioned datasets, Synscapes was created using a parameter-based random generator to build a new scene for each image. Furthermore, the developers leveraged photorealistic rendering techniques and found that high image realism is necessary to bridge the sim-to-real domain gap [Wrenninge.2018]. Nevertheless, the application of autonomous driving is out of scope for industrial manufacturers, so the observed results may not apply to their use cases.
@@
The amount of synthetic image datasets for industrial requirements is small, even though several approaches for image synthesis have been proposed and evaluated in small-scale industrial test cases, as discussed in the following subsection. Two recently published datasets are SIP-17 [Zhu.2023] and SORDI [Akar.2022], aiming to close this gap. SIP-17 is a synthetic dataset for industrial parts classification consisting of 17 objects in six different use cases to allow research in close-to-reality scenarios. The experiments reveal that classification is difficult when classes share a similar albedo or have a comparable shape [Zhu.2023], which is typical for industrial objects.
@@
While SIP-17 contains 33,000 images, SORDI was presented as the first large-scale synthetic image dataset for the industry with 200,000 images at the publication date. It is characterized by eight industrial assets in 32 fully synthetic factory scenarios and enables the sim-to-real benchmarking of vision models for intelligent robots in factory environments [Akar.2022]. However, SORDI cannot serve the entire range of industrial applications, so further work is required to cover, e.g., quality control in assembly and object tracking.
@@
To conclude, it is difficult for industrial manufacturers to estimate how a synthetically trained DL model would perform in a specific use case, as there is not enough data available to benchmark sim-to-real methods in different industry-related applications. Thus, we provide our dataset to enable further investigations in this area with a focus on narrow and very similar industrial objects that are densely arranged and partly occlude each other.
@@ 2.2 Image Synthesis for Deep Learning
Generating synthetic training images for supervised learning tasks requires suitable design choices to improve the sim-to-real generalization performance. These choices include textures of the objects of interest, the image background, different viewpoints, lighting conditions, etc., potentially making the modeling of virtual scenes very time-consuming and complex as well. Especially when the reality should be closely approximated, modeling the entire range of possible real-world variations might eliminate the advantage of saving time for data annotation. Hence, there are various approaches to apply image synthesis that attempt to minimize the domain gap while avoiding excessive modeling effort. For instance, the authors of [Dwibedi.2017] used segmentation masks of real images to cut out the objects of interest and place them onto random background images, which can be seen as image synthesis. This idea was later improved in [Dvornik.2018] proposing a context model that provides patches of the background images likely to contain certain objects to create more realistic training data. Although this method is very efficient and showed promising results in object detection [Dvornik.2018], it relies on real-world data which might not be available or too expensive to gather in industrial applications. Alternatively, the appearance of real-world objects can be approximated utilizing accurate 3D models and modern rendering techniques such as in [Su.2015]. Here, the authors overlaid background images with object renderings, similar to the cut-and-paste approach but completely removing the necessity of annotated real-world data.
@@
Since real images are highly affected by the environment in which they are taken, the appearance of objects changes with their surroundings, impeding model prediction. Hence, a common approach is to randomize the environment of the objects of interest during image synthesis to train the DL model to deal with these variations. However, the authors of [Tobin.2017] hypothesized that complete DR, including even unrealistic variations of the objects of interest, helps to close the sim-to-real domain gap as the DL model is forced to focus on the object’s shape rather than its color or glossiness. Therefore, they also randomized the textures of the 3D models and obtained promising results in the localization of simple geometric objects, enabling a robot to grab these objects in a real-world test [Tobin.2017].
@@
Based on this idea, several experiments have been conducted, assessing different degrees of DR and how domain knowledge can be used during simulation to improve real-world performance. For instance, [Prakash.2019] introduced context splines to the synthetic scenes to ensure that cars are always positioned on streets and [Mayershofer.2021] proposed to define object relations via configuration files to set the objects of interest into more realistic contexts. However, these methods were evaluated on small individual test cases, so the results are hardly comparable and also partly contradicting. This was addressed by the authors of [Eversberg.2021] who compared numerous DR approaches in the application of detecting turbine blades at manual working stations to provide a guideline on designing image synthesis pipelines for industrial use cases. They observed it to be most beneficial to focus on realistic lighting conditions and object textures while the image background and distractors can be randomized completely [Eversberg.2021].
@@
Complementarily, in our research, we follow these guidelines to develop a general strategy for pipeline design in complex industrial parts detection scenarios that ensures a manageable implementation effort and demonstrate the sim-to-real performance of state-of-the-art object detection models in the challenging industrial application of terminal strip object detection, representatively for these scenarios.
@@ 3 Image Synthesis Pipeline
  • The most informative and usually the only accessible view of a terminal strip is the front view, as illustrated in Figure 1.
Since it has been shown to be beneficial to include domain knowledge in DR [Prakash.2019, Mayershofer.2021], we use this information in our pipeline along with standard randomization techniques to obtain random synthetic images that are contextually close to reality. In this section, we provide details about the different generation steps we implemented using the open-source rendering software Blender [^3] Blender 3.3: https://www.blender.org.
@@ 3.1 Creation of Terminal Strips
Apart from wiring, the main distractors of terminal strips are their additional accessories, such as plug-in bridges, test adapters, and markings. Although the authors of [Eversberg.2021] found that realistic distractors do not necessarily improve the performance of DL models, we decided to stay with realism for two reasons. First, the modeling effort is almost identical to simple, unrealistic distractors, as the required 3D models are available, and the virtual materials are similar to the materials used for the terminal blocks. Second, some applications may focus on detecting these accessories, e.g., to determine whether plug-in bridges are plugged correctly. Thus, each terminal strip is randomly equipped with markings, plug-in bridges, and test adapters according to its configuration.
@@ 3.3 Lighting Conditions
In recent image synthesis methods for visual DL applications, there are two common approaches to randomize the lighting conditions: using parameterizable, virtual light sources and implementing image-based lighting. The former is very flexible and allows accurate modeling of the desired illumination. However, creating real-world conditions is laborious since light rays of different colors and intensities might enter the scene from various directions, and reflections should also be considered. In contrast, image-based lighting maps a $360^{\circ}$ image on a globe and uses the pixel values to illuminate the 3D models as if they were in the captured environment. This is less customizable but reduces the creation of complex lighting conditions to the selection of suitable images and has shown to be superior when only little effort is invested in parameterizing virtual light sources [Eversberg.2021].
@@ 3.4 Image Background
Similar to the lighting conditions, synthetic image backgrounds can be virtually created and rendered along with the objects of interest [Akar.2022] or they rely on real images [Eversberg.2021]. On the one hand, modeling the entire background leads to more realistic images, as all objects can be related, and visualizing realistic shadows is possible. On the other hand, random image-based backgrounds have shown promising results [Eversberg.2021] and are less complex, which is more suitable for industrial manufacturers. Nevertheless, the authors of [Alghonaim.2021] found that realistic shadows significantly improve the performance of DL models in a 6D pose estimation task. Therefore, we decided to follow a compromising strategy, reusing the HDRIs of the environment lighting and visualizing the shadow of the terminal strip as if it were mounted on a flat surface.
@@
The implementation is based on a surrounding semi-ellipsoid, as illustrated in Figure 4, that serves as a shadow catcher. It is positioned so that the flat side represents the mounting surface for the DIN rail to which the terminal strip is centered, allowing the objects of interest to cast shadows on the background. To insert the HDRI, the same image from the environment lighting is mapped onto the surface of the semi-ellipsoid, and the respective rotation around the vertical axis is applied. As a result, the background of a synthetic image consists of the ground of the environment captured by the HDRI with additional shadows cast by the terminal strip, which is more realistic than using fully randomized backgrounds and less expensive than modeling the entire scene. See Appendix A for synthetic sample images generated using this approach.